OpenStudy (anonymous):

In triangle ABC, angleA=60 degree, prove that BC2=AB2+AC2+AB.AC

OpenStudy (anonymous):

Can you please double check the question? Is it "-AB.AC" instead of "+AB.AC" ?

OpenStudy (raden):

use the cosine rule :  cos60 = (AB^2 + AC^2 - BC^2)/2ABAC

OpenStudy (raden):

1/2 = (AB^2 + AC^2 - BC^2)/2*AB*AC multiplify by 2 to both side, giving us 1 = (AB^2 + AC^2 - BC^2)/AB*AC

OpenStudy (raden):

cross product out! AB * AC = AB^2 + AC^2 - BC^2

OpenStudy (raden):

and actually, Rolypoly is right, that should "-AB.AC" not +AB AC i think there is a bit typo from u, @kumar69